chocolate drops are $2.30 per pound, gumdrops are $1.50 per pound. You want a total of 50 pounds that is worth $2.10 per pound. How many pounds of each should you use?
You want a total of 50 pounds that is worth $1.00 per pound. Can you do this? Explain.

Answer by erica65404(394)

2.30x+1.50y=2.10*50
x+y=50
y=50-x
2.30x+1.50(50-x)=105
2.30x-1.50x+75=105
.80x=30
x=37.5
50-37.5=y
there are 37.5 lbs chocolate drops and 12.5 ibs gumdrops.
2.30x+1.50y=50
2.30x+1.50(50-x)=50
2.30x-1.50x+75=50
.80x=-25
x=-31.25
It is not possible because you cannot have a negative amount of lbs.